White Phosphorus P consist of P4 molecules it spontaneously ignites on contact with air. If there are 0.350 mole of P4 molecules. how many P4 molecules are present? What is the mass in grams of a sample? How many P atoms are present?
1
Expert's answer
2015-09-29T07:09:09-0400
m(P4) = 0.350 mol * 124g/mol = 43.4 g
N(P4) = 0.350 mol * 6.02 * 10^23 mol^-1 = 2.107 *10^23 molecules
n(P) = 4n(P4)
n(P) = 4*0.350 mol = 1.4 mol;
N(P) = 1.4 mol * 6.02 *10^23 mol^-1 = 8.428*10^23 atoms